Yolanda found the area of Bob's driveway in square meters as shown. Explain why Yolanda's answer is incorrect.
Area = 45 18 = 810 square feet
810 square feet 0.305 meter
_____________
1 foot 247.1 square meters

Answer by Theo(13342)

1 foot = .305 meters.
45 feet * 18 feet = 810 square feet.
45 * .305 * 18 * .305 = 45 * 18 * .305 * .305 = 45 * 18 * .305^2 = 810 * .305^2 = 75.35025 square meters.
she need to multiply square feet by .305^2 to get square meters.